OpenStudy (anonymous):

how do you balance this redox reaction? Cl2+OH^-=ClO3^-+Cl^-+H2O

OpenStudy (anonymous):

I believe this how you do it: For ones like this one of the Cl's in Cl2 is oxidized, one reduced, so write it as Cl + Cl or write it twice Cl2 + Cl2 Cl + Cl = ClO3- + Cl- Cl = ClO3- + 5e- Cl + 1e- = Cl- so Cl + 5Cl = ClO3- + 5Cl- now add the OH- and H2O Cl + 5Cl + 6OH- = ClO3- + 5Cl- + 3H2O now put the Cl's back together 3Cl2 + 6OH- = ClO3- + 5Cl- + 3H2O